# Planetary Edge Trends (PET)

### v1

@@ 4.2 Comparisons with Previous Studies
Mendigutía et al. (2024) combined the Transiting Exoplanet Survey Satellite (TESS) and $Gaia$ DR3 data to study the location of short-period hot Jupiters and found that these planets around intermediate-mass stars tend to orbit closer to their host stars than the inner dust disk, aligning instead with the magnetospheric truncation radius. This suggests that the inner gas disk, rather than the dust disk, constrains the innermost orbits of hot Jupiters in such systems. Based on the above research and comparative analysis, we find that the inner dust disk constrains the innermost orbits of small planets, in contrast to the inner edges of giant planets, which are associated with the magnetospheres of protoplanetary disks (Mendigutía et al., 2024). We believe this comparative result is reasonable and have presented a schematic comparison in Fig. 8. For giant planets, their masses are primarily dominated by gas, and their migration occurs within the gas disk, making the inner edge dependent on the magnetospheric gas-truncation radius. In contrast, small planets are mainly composed of solid material, and their formation relies on the dust disk, resulting in the inner edge being determined by the dust-destruction region. Therefore, the comparison between this study and previous research indicates that the inner edges of different planetary populations (small planets and giant planets) are governed by distinct mechanisms. In summary, the inner edge for giant planets is likely determined by the migration halting mechanism, while for small planets, it is associated with the local solid distribution. This result provides evidence against the hypothesis that most small planets have undergone large-scale inward migration to reach the inner edge.
@@
Plavchan & Bilinski (2013) used short-period confirmed Jovian exoplanets and $Kepler$ planetary candidates from the early data release to study the relationship between the semi-major axis of all planets and stellar mass (planet frequency as a function of stellar mass). They found that the tidal halting mechanism provided the best fit for confirmed Jovian exoplanets. However, for $Kepler$ candidates, they did not confirm a specific mechanism but noted that the power-law index $\gamma_{1}=0.38-0.9$ is larger than the predicted range for the tidal halting mechanism ($0.23-0.33$). In contrast, by using only the innermost planet in multis as the inner edge to study its relationship with stellar mass, we find a power-law index of $\gamma_{1}=0.6-1.1$. This result is comparable to Plavchan & Bilinski (2013)’s finding for $Kepler$ candidates, although the samples are not entirely the same.
@@
Mulders et al. (2015) studied sub-Neptunes and super-Earths ($1\,R_{\oplus}$ to $4\,R_{\oplus}$) to investigate the occurrence of small planets around stars of different masses. They found an occurrence drop at an orbital period of approximately 10 days, regardless of stellar mass. Using this occurrence drop to trace the inner edge of planetary systems, they derived a relationship of $a_{\text{in}}\propto M_{\star}^{1/3}$, suggesting that the co-rotation radius best matches the stellar-mass-dependent location of the inner edge. However, by using the innermost planets in multis as the inner edge to study the relationship with stellar mass, we find a stronger correlation with stellar mass (around $a_{\text{in}}\propto M_{\star}^{0.6-1.1}$), which aligns well with the prediction of dust sublimation model. We find that the discrepancy stems from using planet occurrence rates to trace the inner edge, which underestimates the correlation index ($\gamma_{1}$) and diminishes the influence of stellar mass on the location of inner-edge planets. A more detailed discussion and analysis is provided in Appendix A. Previous studies analyze the position of the inner edge using planet occurrence rates (such as Eqs. (7), (8), and (9) in Mulders et al., 2015), which consider all planets in the system. In this method, outer planets will have more weight (due to correction of transit geometric effect) to determine the inner edge than the innermost one, although the innermost one represents the true inner edge. Consequently, in multis, lower-mass stars tend to host more planets (Yang et al., 2020), shifting the inner edge derived from occurrence rates outward. This shift results in a weaker correlation between the inner edge and stellar mass compared to the intrinsic correlation. As shown in Appendix A, if the intrinsic correlation between the inner edge and stellar mass has a power-law index of 1.0, it would be underestimated as 0.3 when using the occurrence rate method. This effectively explains the discrepancy between the results of Mulders et al. (2015) and this work.
@@
Additionally, Howard et al. (2012), Mulders et al. (2015), Yang et al. (2020), and Giacalone & Dressing (2025) have all found that close-in planets are rarer around higher-mass stars, a result qualitatively consistent with that as stellar mass increases, the position of the inner edge shifts outward. Compared to FGKM-type stars, A-type stars have significantly higher temperatures (with more pronounced differences in temperature than stellar mass). As a result, the dust sublimation radius shifts outward, pushing the inner edge farther away and leaving little room for small, close-in planets to form. Using the relationship between the inner edge and stellar mass derived in this work (All-MLR-$\gamma_{1}$: $0.812_{-0.078}^{+0.085}$), we extrapolate to A-type stars ($\sim 2\,M_{\odot}$) and find that the inner edge would reach $0.21-0.27$ AU. This implies that it is challenging for planets to exist within 0.2 AU around A-type stars, which aligns well with the results of Giacalone & Dressing (2025), demonstrating that close-in planets orbiting A-type stars are rare.
@@ 4.3 Inspirations for Future Work
Although the observational results generally align with some of the theoretical models in Fig. 7, we can see that there is a significant scatter in the observational data, reflected in large uncertainties in the intercepts. The intercept essentially acts as a normalization factor (e.g., $\log(\gamma_{0})$ in Eqs. (3) and (8)). This may be related to two factors: the surface density of the protoplanetary disk and the size of the grains or dust within the disk (see Fig. 8 in Kama et al., 2009). Thus, the large scatter of observational data around theoretical models (Fig. 7) may reflect the diversity in the initial conditions of the protoplanetary disk. Furthermore, the relationship between the inner edge and stellar mass may be influenced by more complex factors. For instance, some studies (Liu et al., 2017; Liu & Ormel, 2017) suggest that after planet formation, the disk dissipation process further affects the position of the inner edge, as the orbits of planets expand with the inner magnetospheric cavity’s growth. In addition, other dynamical processes at later stages, such as giant impacts (Izidoro et al., 2017; Gabriel & Cambioni, 2023), planetary scattering (Ford & Rasio, 2008; Chatterjee et al., 2008), and secular chaos (Laskar, 1997; Lithwick & Wu, 2011), may further shape the architecture of planetary systems, changing the position of the innermost planets. Therefore, future theoretical model simulations should take into account more factors, as the strength of the scatter is influenced by multiple elements. Exploring the combinations of these factors and their combined effects would be a direction for further research.
@@
In this work, focusing on small planets, we find that different populations exhibit varying correlations between the inner edge and stellar mass (Table 2). Among them, mixed multiple systems show the strongest correlation, while super-Earths and sub-Neptunes exhibit much weaker correlations. This may be because small planets are inherently diverse and complex, with different populations undergoing distinct formation and evolution processes that may correspond to different theoretical mechanisms. However, the sample size in this work is still relatively small. Future studies with more data are needed to further validate and explore these findings.
@@
Apart from stellar mass, which is the primary focus of this work, the inner edge is likely determined by multiple properties. The correlation between the inner edge and stellar mass may reflect the role of dust sublimation in shaping the inner edge. However, the theoretical understanding of how other properties, such as stellar metallicity, influence the inner edge remains unclear. This could also contribute to the significant scatter observed. In addition, stellar age is another important property. Planetary systems evolve over time, and stellar age is also correlated with other factors. Therefore, analyzing the influence of stellar age is crucial for understanding the long-term evolution of planetary systems. In the future, further theoretical and observational work is needed to fully understand the influence of properties such as stellar metallicity and stellar age on the inner edge.
@@
The edges of planetary systems are essential for understanding the mechanisms of planet formation. While this work focuses on the inner edge, it is important to note that the outer edge is equally significant. Millholland et al. (2022) and Sobski & Millholland (2023), in their studies of the outer edges of $Kepler$ multis, conducted detailed analyses and demonstrated that the existence of the outer edge is not a result of observational limitations, but rather due to specific physical mechanisms. Nonetheless, the theoretical truncation mechanism responsible for the outer edge remains unknown. Therefore, we are curious to explore whether there is any connection between the inner edge and the outer edge. There are still numerous intriguing questions that warrant further exploration in our future work.
@@ 5 Conclusion and Summary
The position of the innermost planet in a planetary system reflects the relationship between the entire system and its host star, offering potentially significant insights into planetary formation and evolution processes. However, previous studies have revealed persistent discrepancies between observational results and theoretical models, leaving the primary mechanism governing the position of the inner edge in planetary systems uncertain. In this study, we use the $Kepler$ DR25 catalog, combined with LAMOST and $Gaia$ data, focusing specifically on small planets in multis (Table 1 and Fig. 1). The data are categorized into four planetary populations (Fig. 2): All Multiple Systems, Super-Earth Systems, Sub-Neptune Systems, and Mixed Multiple Systems, aiming to study the correlation between stellar mass and the position of the inner edge.
Our findings reveal that as stellar mass increases, the position of the inner edge also increases (Fig. 3). Simultaneously, we observe that the inner edges of hot small-planet systems decrease with stellar metallicity (Fig. 4). Since stellar mass is also correlated with metallicity, therefore the intrinsic correlation between the inner edge and stellar mass should be even stronger. After correcting for metallicity dependence using a Multiple Linear Regression Model, we observe a steeper correlation between stellar mass and the inner edge (Fig. 5). Analyzing the dataset across four planetary populations reveals a correlation power-law index of $\gamma_{1}=0.6-1.1$. In addition, we find that observational biases do not affect our results (Fig. 6).
In analyzing the observational data, we identify that using occurrence rates, as employed in previous studies, introduces a bias that underestimates the dependence of inner-edge planets on stellar mass (Appendix A). Correcting for this influence reveals a stronger relationship between the inner edge and stellar mass. This effect explains the discrepancy between this work ($\gamma_{1}=0.6-1.1$) and a previous study ($\gamma_{1}=0.3$, Mulders et al., 2015).
By comparing our results with existing theoretical models (Table 2 and Fig. 7), we find that the observed correlation between the inner edge and stellar mass aligns with the pre-main-sequence (PMS) dust sublimation radius of the protoplanetary disk. This supports the hypothesis that dust sublimation is an important mechanism that affects the formation and orbital evolution of small planets (Boley & Ford, 2013; Flock et al., 2016), contrasting with the inner edges of hot Jupiters, which are influenced by the magnetospheres of protoplanetary disks (Mendigutía et al., 2024). This underscores that the inner edge of different planetary populations may indeed be governed by distinct mechanisms (Fig. 8).
@@ Acknowledgements.
We thank Qi Hao, Jiayi Yang, and Shuting Han for their valuable discussions on methods, analyses, and figures. We also extend our gratitude to the referee for their insightful comments.
@@ Appendix A Occurrence Rates Underestimate the Inner Edge Stellar Mass Correlation
Figure 9: A schematic diagram illustrating the impact of planetary occurrence rates on the conclusion. In the diagram, $M^{\text{L}}$ and $M^{\text{H}}$ represent lower-mass and higher-mass host stars, respectively, while $N^{\text{L}}$ and $N^{\text{H}}$ denote the number of planets in each system ($N^{\text{L}}>N^{\text{H}}$). $a_{\text{in}}^{\text{L}}$ and $a_{\text{in}}^{\text{H}}$ represent the semi-major axes of the innermost planets in the two systems. $c^{\text{L}}$ and $c^{\text{H}}$ represent the conversion factor between the period ratio and the semi-major axis ratio of the planetary distribution in the system.
@@
We further conduct a quantitative analysis to specifically calculate the extent to which the relationship between the inner edge and stellar mass has been weakened. According to equations from Mulders et al. (2015), the occurrence rate for a planet of a given radius and orbital period is:
@@
$f_{\text{geo}}=\frac{R_{\text{p}}+R_{\star}}{a(1-e^{2})},$ (13)
$N_{\star}(\{T_{\text{eff}}\},R_{\text{p}},P)=\sum_{i=0}^{N_{\star}(\{T_{\text{eff}}\})}\left(f_{\text{eff},\text{i}}\cdot f_{\text{n},\text{i}}\right),$ (14)
and round $N_{\star}$ to the nearest integer. Considering that the $Kepler$ system we are studying consists of nearly circular orbits ($e=0$) and $R_{\text{p}}\ll R_{\star}$, for a given system, $R_{\star}$ and $N_{\star}$ remain consistent. Therefore, we can derive that $f_{\text{occ}}\propto a_{\text{occ}}$. Following Mulders et al. (2015), the inner edge calculated from occurrence rates ($a_{\text{occ}}$) is defined at the point where $f_{\text{occ}}$ drops by half (or the mean point). Thus, we can derive:
@@
where $a_{\text{occ}}$ represents the inner edge of a planetary system calculated through $f_{\text{occ}}$, $N$ represents the number of planets in the system, and $a_{\text{n}}$ denotes the semi-major axis of each planet in the system, ordered from innermost to outermost. We assume that the semi-major axis ratio between two adjacent planets in a system is constant and related to their period ratio, which can be calculated through $a=\sqrt[3]{GM_{\star}P^{2}/4\pi^{2}}$. Thus, we can further derive:
@@
where $a_{\text{in}}$ represents the real position of the innermost planet, and $c=a_{\text{n}+1}/a_{\text{n}}=\left(P_{\text{n}+1}/P_{\text{n}}\right)^{2/3}$ represents the conversion process between the period ratio and the semi-major axis ratio for planets in the system.
To facilitate quantitative calculations and to visually observe the impact of planet occurrence rates, we hypothesize two planetary systems, as shown in Fig. 9. $M^{\text{L}}$ and $M^{\text{H}}$ represent lower-mass and higher-mass host stars, respectively, while $N^{\text{L}}$ and $N^{\text{H}}$ denote the number of planets in each system ($N^{\text{L}}>N^{\text{H}}$). $a_{\text{in}}^{\text{L}}$ and $a_{\text{in}}^{\text{H}}$ represent the semi-major axes of the innermost planets in the two systems. $c^{\text{L}}$ and $c^{\text{H}}$ represent the semi-major axis ratios of adjacent planets in two respective systems. For these two systems, we can derive the following formulas:
$f_{\text{occ}}^{\text{L}}\propto a_{\text{occ}}^{\text{L}}=\frac{a_{\text{in}}^{\text{L}}\cdot\sum_{\text{n}=0}^{N^{\text{L}}-1}(c^{\text{L}})^{\text{n}}}{N^{\text{L}}},$ (17)
and
$f_{\text{occ}}^{\text{H}}\propto a_{\text{occ}}^{\text{H}}=\frac{a_{\text{in}}^{\text{H}}\cdot\sum_{\text{n}=0}^{N^{\text{H}}-1}(c^{\text{H}})^{\text{n}}}{N^{\text{H}}}.$ (18)
@@
In this work, we express the correlation between the inner edge and stellar mass using $a_{\text{in}}\propto M_{\star}^{\gamma_{1}}$ and $a_{\text{occ}}\propto M_{\star}^{\gamma_{\text{occ}}}$, then we have:
$\frac{a_{\text{occ}}^{\text{H}}}{a_{\text{occ}}^{\text{L}}}=(\frac{M^{\text{H}}}{M^{\text{L}}})^{\gamma_{\text{occ}}},$ (20)
and
$\frac{a_{\text{in}}^{\text{H}}}{a_{\text{in}}^{\text{L}}}=(\frac{M^{\text{H}}}{M^{\text{L}}})^{\gamma_{1}}.$ (21)
Thus, through Eqs. (20) and (21), Eq. (19) can be transformed into:
@@
Figure 10: The period ratios of all the systems studied in this work. From left to right, the host star mass of the planetary systems gradually increases.

### v2

@@ 4.2 Comparisons with previous studies
Mendigutía et al. (2024) combined the Transiting Exoplanet Survey Satellite (TESS) and $Gaia$ DR3 data to study the location of short-period hot Jupiters and found that these planets around intermediate-mass stars tend to orbit closer to their host stars than the inner dust disk, aligning instead with the magnetospheric truncation radius. This suggests that the inner gas disk, rather than the dust disk, constrains the innermost orbits of hot Jupiters in such systems. Based on the above research and comparative analysis, we find that the inner dust disk constrains the innermost orbits of small planets, in contrast to the inner edges of giant planets, which are associated with the magnetospheres of protoplanetary disks (Mendigutía et al., 2024). It is important to note that their study employs a different research methodology from ours. Different objectives require different analytical approaches. Here, we primarily conduct a preliminary discussion and comparison of the conclusions drawn from the two studies. We believe this comparative result is reasonable and have presented a schematic comparison in Fig. 8. For giant planets, their masses are primarily dominated by gas, and their migration occurs within the gas disk, making the inner edge dependent on the magnetospheric gas-truncation radius. In contrast, small planets are mainly composed of solid material, and their formation relies on the dust disk, resulting in the inner edge being determined by the dust-destruction region. Therefore, the comparison between this study and previous research indicates that the inner edges of different planetary populations (small planets and giant planets) are governed by distinct mechanisms. In summary, the inner edge for giant planets is likely determined by the migration halting mechanism, while for small planets, it is associated with the local solid distribution.
Figure 8: Schematic comparison of the conclusions of this work and Mendigutía et al. (2024), revealing that the inner edges of different planetary systems correspond to different mechanisms. The results obtained in this work for small planets align more closely with the dust sublimation model, whereas their results obtained for giant planets are more consistent with migration up to the gas-truncation radius.
While the alignment between the inner disk edge and the dust sublimation radius diminishes the necessity of orbital migration for small planets, migration cannot be entirely discounted provided that it ceases effectively near the sublimation boundary. For example, Zhu et al. (2024) proposed that the inner dead zone edge, rather than the magnetospheric truncation radius, determined a planet’s final position before the protoplanetary disc dissipated. As illustrated in Fig. 23 of their work, this effect appeared to be more pronounced for low-mass planets. While high-mass planets might still have migrated inward beyond the dust sublimation radius, low-mass planets were suggested to be unable to migrate within the inner magneto-rotational instability (MRI)-active region. Instead, they were likely to be trapped at the dead zone, which nearly coincided with the dust sublimation radius—where dust sublimated and MRI could operate. Thus, our observational results provide support for their theoretical simulation model, as the findings appear to be broadly consistent with our observations.
@@
Plavchan & Bilinski (2013) used short-period confirmed Jovian exoplanets and $Kepler$ planetary candidates from the early data release to study the relationship between the semimajor axis of all planets and stellar mass (planet frequency as a function of stellar mass). They found that the tidal halting mechanism provided the best fit for confirmed Jovian exoplanets. However, for $Kepler$ candidates, they did not confirm a specific mechanism but noted that the power-law index $\gamma_{1}=0.38-0.9$ is larger than the predicted range for the tidal halting mechanism ($0.23-0.33$). In contrast, by using only the innermost planet in multis as the inner edge to study its relationship with stellar mass, we find a power-law index of $\gamma_{1}=0.6-1.1$. This result is comparable to Plavchan & Bilinski (2013)’s finding for $Kepler$ candidates, although the samples are not entirely the same.
@@
Mulders et al. (2015) studied sub-Neptunes and super-Earths ($1\,R_{\oplus}$ to $4\,R_{\oplus}$) to investigate the occurrence of small planets around stars of different masses. They found an occurrence drop at an orbital period of approximately 10 days, regardless of stellar mass. Using this occurrence drop to trace the inner edge of planetary systems, they derived a relationship of $a_{\text{in}}\propto M_{\star}^{1/3}$, suggesting that the co-rotation radius best matches the stellar-mass-dependent location of the inner edge. However, by using the innermost planets in multis as the inner edge to study the relationship with stellar mass, we find a stronger correlation with stellar mass (around $a_{\text{in}}\propto M_{\star}^{0.6-1.1}$), which aligns well with the prediction of dust sublimation model. We find that the discrepancy stems from using planet occurrence rates to trace the inner edge, which underestimates the correlation index ($\gamma_{1}$) and diminishes the influence of stellar mass on the location of inner-edge planets. A more detailed discussion and analysis is provided in appendix A. Previous studies analyze the position of the inner edge using planet occurrence rates (such as Eqs. (7), (8), and (9) in Mulders et al., 2015), which consider all planets in the system. In this method, outer planets will have more weight (due to the correction of the transit geometric effect) to determine the inner edge than the innermost one, although the innermost one represents the true inner edge. Consequently, in multis, lower-mass stars tend to host more planets (Yang et al., 2020), shifting the inner edge derived from occurrence rates outward. This shift results in a weaker correlation between the inner edge and stellar mass compared to the intrinsic correlation. As shown in appendix A, if the intrinsic correlation between the inner edge and stellar mass has a power-law index of 1.0, it would be underestimated as 0.3 when using the occurrence rate method. This effectively explains the discrepancy between the results of Mulders et al. (2015) and this work.
@@
Additionally, Howard et al. (2012), Mulders et al. (2015), Yang et al. (2020), and Giacalone & Dressing (2025) have all found that close-in planets are rarer around higher-mass stars, a result qualitatively consistent with that as stellar mass increases, the position of the inner edge shifts outward. Compared to FGKM-type stars, A-type stars have significantly higher temperatures (with more pronounced differences in temperature than stellar mass). As a result, the dust sublimation radius shifts outward, pushing the inner edge farther away and leaving little room for small, close-in planets to form. Using the relationship between the inner edge and stellar mass derived in this work (All-MLR-$\gamma_{1}$: $0.81_{-0.08}^{+0.09}$), we extrapolate to A-type stars ($\sim 2\,M_{\odot}$) and find that the inner edge would reach $0.21-0.27$ AU. This implies that it is challenging for planets to exist within 0.2 AU around A-type stars, which aligns well with the results of Giacalone & Dressing (2025), demonstrating that close-in planets orbiting A-type stars are rare.
@@ 4.3 Inspirations for future work
Although the observational results generally align with some of the theoretical models in Fig. 7, we can see that there is a significant scatter in the observational data, reflected in large uncertainties in the intercepts. The intercept essentially acts as a normalization factor (e.g., $\log(\gamma_{0,am})$ in Eq. (3) and $\log(\gamma_{0,amf})$ in Eq. (8)). This may be related to two factors: the surface density of the protoplanetary disk and the size of the grains or dust within the disk (see Fig. 8 in Kama et al., 2009). Thus, the large scatter of observational data around theoretical models (Fig. 7) may reflect the diversity in the initial conditions of the protoplanetary disk. Furthermore, the relationship between the inner edge and stellar mass may be influenced by more complex factors. For instance, some studies (Liu et al., 2017; Liu & Ormel, 2017) suggest that after planet formation, the disk dissipation process further affects the position of the inner edge, as the orbits of planets expand with the inner magnetospheric cavity’s growth. In addition, other dynamical processes at later stages, such as giant impacts (Izidoro et al., 2017; Gabriel & Cambioni, 2023), planetary scattering (Ford & Rasio, 2008; Chatterjee et al., 2008), and secular chaos (Laskar, 1997; Lithwick & Wu, 2011), may further shape the architecture of planetary systems, changing the position of the innermost planets. Future work could distinguish between resonant and nonresonant systems to further explore the relationship between the inner edge and stellar parameters. In this context, we emphasize the potential role of evolutionary processes. However, studying such effects requires precise stellar age measurements, which remain challenging due to large uncertainties in current data. Despite these limitations, understanding the evolutionary history of planetary systems remains an important goal. Therefore, future theoretical model simulations should take into account more factors, as the strength of the scatter is influenced by multiple elements. Exploring the combinations of these factors and their combined effects would be a direction for further research.
@@
In this work, focusing on small planets, we find that different populations exhibit varying correlations between the inner edge and stellar mass (Table 2). Among them, mixed multiple systems show the strongest correlation, while super-Earths and sub-Neptunes exhibit much weaker correlations. This may be because small planets are inherently diverse and complex, with different populations undergoing distinct formation and evolution processes that may correspond to different theoretical mechanisms. However, the sample size in this work is still relatively small. Future studies with more data are needed to further validate and explore these findings. Additionally, to assess the impact of radius uncertainties on planet classification, we conduct simulations in which planets are repeatedly reclassified and refitted. The resulting slope distributions for super-Earths and sub-Neptunes are $0.57_{-0.10}^{+0.10}$ and $0.45_{-0.17}^{+0.17}$, respectively. These results show increased uncertainties compared to the observed values of $0.54_{-0.05}^{+0.05}$ and $0.38_{-0.09}^{+0.09}$, which remain consistent within $1\sigma$. Thus, incorporating radius uncertainties enlarges the error bars and reduces the statistical significance of the difference between the two populations. Given the small sample sizes, this difference remains suggestive rather than conclusive.
@@
Apart from stellar mass, which is the primary focus of this work, the inner edge is likely determined by multiple properties. The correlation between the inner edge and stellar mass may reflect the role of dust sublimation in shaping the inner edge. However, the theoretical understanding of how other properties, such as stellar metallicity, influence the inner edge remains unclear. This could also contribute to the significant scatter observed. In addition, stellar age is another important property. Planetary systems evolve over time, and stellar age is also correlated with other factors. Therefore, analyzing the influence of stellar age is crucial for understanding the long-term evolution of planetary systems. In the future, further theoretical and observational work is needed to fully understand the influence of properties such as stellar metallicity and stellar age on the inner edge. In addition, larger planets also represent an interesting research direction. However, since this study already covers an extensive scope and categorizes the sample into super-Earths and sub-Neptunes, we do not further explore the impact of larger planets in this work. Nonetheless, we suggest that future research incorporate planetary mass or radius into the relationship between the inner edge and stellar mass for a more comprehensive analysis.
@@
The edges of planetary systems are essential for understanding the mechanisms of planet formation. While this work focuses on the inner edge, it is important to note that the outer edge is equally significant. Millholland et al. (2022) and Sobski & Millholland (2023), in their studies of the outer edges of $Kepler$ multis, conducted detailed analyses and demonstrated that the existence of the outer edge is not a result of observational limitations, but rather due to specific physical mechanisms. Nonetheless, the theoretical truncation mechanism responsible for the outer edge remains unknown. Therefore, we are curious to explore whether there is any connection between the inner edge and the outer edge.
@@ 5 Conclusion and summary
The position of the innermost planet in a planetary system reflects the relationship between the entire system and its host star, offering potentially significant insights into planetary formation and evolution processes. However, previous studies have revealed persistent discrepancies between observational results and theoretical models, leaving the primary mechanism governing the position of the inner edge in planetary systems uncertain. In this work, we used the $Kepler$ DR25 catalog combined with LAMOST and $Gaia$ data and focused specifically on small planets in multis (Table 1 and Fig. 1). The data were categorized into four planetary populations (Fig. 2): all multiple systems, super-Earth systems, sub-Neptune systems, and mixed multiple systems. Our aim was to study the correlation between stellar mass and the position of the inner edge.
Our findings revealed that as stellar mass increases, the position of the inner edge also increases (Fig. 3). Simultaneously, we observed that the inner edges of hot small-planet systems decrease with stellar metallicity (Fig. 4). Since stellar mass is also correlated with metallicity, the intrinsic correlation between the inner edge and stellar mass should therefore be even stronger. After correcting for metallicity dependence using an MLR model, we observed a steeper correlation between stellar mass and the inner edge (Fig. 5). Analyzing the dataset across four planetary populations revealed a correlation power-law index of $\gamma_{1}=0.6-1.1$. In addition, we found that observational biases do not affect our results (Fig. 6).
In analyzing the observational data, we identified that using occurrence rates, as employed in previous studies, introduces a bias that underestimates the dependence of inner-edge planets on stellar mass (appendix A). Correcting for this influence reveals a stronger relationship between the inner edge and stellar mass. This effect explains the discrepancy between this work ($\gamma_{1}=0.6-1.1$) and a previous study ($\gamma_{1}=0.3$, Mulders et al., 2015).
By comparing our results with existing theoretical models (Table 2 and Fig. 7), we found that the observed correlation between the inner edge and stellar mass aligns with the pre-main-sequence dust sublimation radius of the protoplanetary disk. This supports the hypothesis that dust sublimation is an important mechanism that affects the formation and orbital evolution of small planets (Boley & Ford, 2013; Flock et al., 2016), contrasting with the inner edges of hot Jupiters, which are influenced by the magnetospheres of protoplanetary disks (Mendigutía et al., 2024). This underscores the possibility that the inner edge of different planetary populations may indeed be governed by distinct mechanisms (Fig. 8).
@@ Acknowledgements.
We thank Qi Hao, Jiayi Yang, and Shuting Han for their valuable discussions on methods, analyses, and figures. We also extend our gratitude to the referee for their insightful comments. This work is supported by the National Key R&D Program of China (No. 2024YFA1611803) and the National Natural Science Foundation of China (NSFC; Grant Nos. 12273011, 12150009, 12222303, and 12173035). J.-W. X. also acknowledges the support from the National Youth Talent Support Program. B. L. is also supported by the start-up grant of the Bairen program from Zhejiang University. Funding for LAMOST (www.lamost.org) has been provided by the Chinese NDRC. LAMOST is operated and managed by the National Astronomical Observatories, CAS.
@@ Appendix A Occurrence rates underestimate the inner edge stellar mass correlation
Figure 9: Schematic diagram illustrating the impact of planetary occurrence rates on the conclusion. In the diagram, $M^{\text{L}}$ and $M^{\text{H}}$ represent lower-mass and higher-mass host stars, respectively, while $N^{\text{L}}$ and $N^{\text{H}}$ denote the number of planets in each system ($N^{\text{L}}>N^{\text{H}}$). $a_{\text{in}}^{\text{L}}$ and $a_{\text{in}}^{\text{H}}$ represent the semimajor axes of the innermost planets in the two systems. $c^{\text{L}}$ and $c^{\text{H}}$ represent the conversion factor between the period ratio and the semimajor axis ratio of the planetary distribution in the system.
We further conduct a quantitative analysis to specifically calculate the extent to which the relationship between the inner edge and stellar mass has been weakened. According to equations from Mulders et al. (2015), the occurrence rate for a planet of a given radius and orbital period is
@@
$f_{\text{geo}}=\frac{R_{\text{p}}+R_{\star}}{a(1-e^{2})}\,,\;N_{\star}(\{T_{\text{eff}}\},R_{\text{p}},P)=\sum_{i=0}^{N_{\star}(\{T_{\text{eff}}\})}\left(f_{\text{eff},\text{i}}\cdot f_{\text{n},\text{i}}\right)\,,\;$ (13)
and $N_{\star}$ is rounded to the nearest integer. Considering that the $Kepler$ system we are studying consists of nearly circular orbits ($e=0$) and $R_{\text{p}}\ll R_{\star}$, for a given system, $R_{\star}$ and $N_{\star}$ remain consistent. Therefore, we can derive that $f_{\text{occ}}\propto a_{\text{occ}}$. Following Mulders et al. (2015), the inner edge calculated from occurrence rates ($a_{\text{occ}}$) is defined at the point where $f_{\text{occ}}$ drops by half (or the mean point). Thus, we can derive
@@
where $a_{\text{occ}}$ represents the inner edge of a planetary system calculated through $f_{\text{occ}}$, $N$ represents the number of planets in the system, and $a_{\text{n}}$ denotes the semimajor axis of each planet in the system, ordered from innermost to outermost. We assume that the semimajor axis ratio between two adjacent planets in a system is constant and related to their period ratio, which can be calculated through $a=\sqrt[3]{GM_{\star}P^{2}/4\pi^{2}}$. Thus, we can further derive
@@
where $a_{\text{in}}$ represents the real position of the innermost planet, and $c=a_{\text{n}+1}/a_{\text{n}}=\left(P_{\text{n}+1}/P_{\text{n}}\right)^{2/3}$ represents the conversion process between the period ratio and the semimajor axis ratio for planets in the system.
To facilitate quantitative calculations and to visually observe the impact of planet occurrence rates, we hypothesize two planetary systems, as shown in Fig. 9. $M^{\text{L}}$ and $M^{\text{H}}$ represent lower-mass and higher-mass host stars, respectively, while $N^{\text{L}}$ and $N^{\text{H}}$ denote the number of planets in each system ($N^{\text{L}}>N^{\text{H}}$). $a_{\text{in}}^{\text{L}}$ and $a_{\text{in}}^{\text{H}}$ represent the semimajor axes of the innermost planets in the two systems. $c^{\text{L}}$ and $c^{\text{H}}$ represent the semimajor axis ratios of adjacent planets in two respective systems. For these two systems, we can derive the following formulas:
$f_{\text{occ}}^{\text{L}}\propto a_{\text{occ}}^{\text{L}}=\frac{a_{\text{in}}^{\text{L}}\cdot\sum_{\text{n}=0}^{N^{\text{L}}-1}(c^{\text{L}})^{\text{n}}}{N^{\text{L}}}\,,\;f_{\text{occ}}^{\text{H}}\propto a_{\text{occ}}^{\text{H}}=\frac{a_{\text{in}}^{\text{H}}\cdot\sum_{\text{n}=0}^{N^{\text{H}}-1}(c^{\text{H}})^{\text{n}}}{N^{\text{H}}}.$ (16)
@@
In this work, we express the correlation between the inner edge and stellar mass using $a_{\text{in}}\propto M_{\star}^{\gamma_{1}}$ and $a_{\text{occ}}\propto M_{\star}^{\gamma_{\text{occ}}}$; then we have
$\frac{a_{\text{occ}}^{\text{H}}}{a_{\text{occ}}^{\text{L}}}=(\frac{M^{\text{H}}}{M^{\text{L}}})^{\gamma_{\text{occ}}}\,,\;\frac{a_{\text{in}}^{\text{H}}}{a_{\text{in}}^{\text{L}}}=(\frac{M^{\text{H}}}{M^{\text{L}}})^{\gamma_{1}}.$ (18)
Thus, through Eq. (18), Eq. (17) can be transformed into
@@
Figure 10: Period ratios of all the systems studied in this work. From left to right, the host star mass of the planetary systems gradually increases.